# Vector-Valued Holomorphic Functions and Abstract Fubini-Type Theorems

### v1

@@ Abstract.
vector-valued, holomorphic functions of several variables
@@ 1. Introduction
The paper is organized as follows. In Section 2 we recall classical results for vector-valued holomorphic functions. Next, in Section 3, we consider the case of a closed subspace $E$ of the space $\mathrm{BM}(\Omega,\Sigma)$ of bounded measurable functions. The main result of that section is Theorem 3.1, which is new (as far as we know). The second main result, Theorem 4.2, is presented in Section 4 and treats the case $E=\mathrm{L}_{p}(\Omega)$ for $1\leq p<\infty$. It is complemented by Theorem 4.3, which deals with $E=\mathrm{L}_{\infty}(\Omega)$.
@@
A subset $M\subseteq\ell^{\infty}(\Omega;E)$ is called (sequentially) bp-closed if it is closed under taking bp-limits of sequences in $M$.
Let $F\subseteq\ell^{\infty}(\Omega;E)$ be a bp-closed subspace. A linear mapping $T:F\to X$ (where $X$ is any topological vector space) is called (sequentially) bp-continuous if
@@
For any measurable space $(\Omega,\Sigma)$ we let $\mathrm{BM}(\Omega):=\mathrm{BM}(\Omega,\Sigma)$ be the Banach space of all bounded and measurable $\mathbb{C}$-valued functions, endowed with the supremum norm. This is a (sequentially) bp-closed subspace of $\ell^{\infty}(\Omega)$.
@@ 2. Vector-Valued Multivariate Holomorphic Functions
Here and in the following, $O\subseteq\mathbb{C}^{d}$ is a fixed open and non-empty set and $E$ is a complex Banach space. A function $f:O\to E$ is holomorphic if it is totally differentiable (= Fréchet-differentiable) with $\mathbb{C}$-linear derivative at each point of $O$ [11, Def. 147, p.68]. The following theorem is a useful characterization of holomorphy. It extends well-known results for multi-variable scalar-valued and one-variable vector-valued functions.
@@
We first show as in [13, proof of 1.3] that $f$ is continuous at $a$. Let $z=(z_{j})_{j}\in D$ and write
@@
is a well-defined $E$-valued holomorphic function on $D$. Indeed, $g$ is certainly holomorphic in each variable separately and each partial derivative is continuous. Hence one can apply [14, XIII, Thm. 7.1].
@@
1) For univariate functions, Theorem 2.1 is well-known, see [3, Appendix A]. For scalar-valued functions, the equivalence (i) $\Leftrightarrow$ (ii) is called Osgood’s lemma; the (a priori) stronger equivalence (i) $\Leftrightarrow$ (iii) is [13, Thm. 1.3].
2) Most books start from a different definition of holomorphy than ours and do not even mention Fréchet-differentiability. A noteworthy exception is [11]. See also [11, Theorem 160] for further characterizations of holomorphy.
3) Assertion (iii) involving a norming subset is due to Grothendieck [8], see also [12, p.139]. It implies (via the Hahn–Banach theorem) the following equivalences:
@@
    A function $f:O\to E$ is holomorphic if and only if it is weakly holomorphic (Dunford’s theorem, see [11, Thm. 148, p.68]).
@@
    A function $T:O\to\mathcal{L}(E;F)$ is holomorphic if and only if it is strongly holomorphic, if and only if it is locally bounded and for each $x\in E$ and each $x^{\prime}\in N$ from a norming subset $N\subseteq F^{\prime}$ the function $\langle F(\cdot)x,x^{\prime}\rangle$ is holomorphic.
4) In (iii) one can weaken the assumptions and require merely that $N$ is separating. This follows from the analogous result for univariate functions due to Arendt and Nikolski [4, Thm. 3.1], see also [2].
@@
This follows from Hartogs’ theorem, which says that a scalar multi-variable function is already holomorphic if it is merely separately holomorphic [11, Thm. 153, p.69].
@@
We write $\mathrm{H}^{\infty}(O;E)$ for the space of all bounded holomorphic $E$-valued functions on $O$.
If we endow $O$ with the Borel $\sigma$-algebra, $\mathrm{H}^{\infty}(O;E)$ becomes a (sequentially) bp-closed subspace of $\mathrm{BM}(O;E)$. From the Cauchy integral formula it follows that for fixed $a\in O$ and $\alpha\in\mathbb{N}_{0}^{d}$ the mapping
@@
3. Holomorphic Families of Bounded Functions
Our first abstract Fubini-type theorem involves closed subspaces of bounded functions. Note that every space $E=\mathrm{BM}(\Omega,\Sigma)$ for some $\sigma$-algebra $\Sigma$ on a set $\Omega$ is a closed subspace of $\ell^{\infty}(\Omega)$.
@@ 3. Holomorphic Families of Bounded Functions
Note that the second part of b) is of the type “differentiation under the integral” when one is inclined to interpret the application of $\mu$ as some kind of integral. We shall break up the proof into two parts.
Proof of Theorem 3.1, a) and b).
@@
are all holomorphic, $F$ is holomorphic by Theorem 2.1.
It follows that $D^{\alpha}F$ takes values in $E$ and
@@
as scalar functions on $O$ for each $\mu\in E^{\prime}$ (Lemma 2.4). Specializing $\mu=\delta_{t}$ for $t\in\Omega$ yields $(D^{\alpha}F)(z)(t)=D^{\alpha}_{z}f(z,t)$, and hence b) is proved. ∎
We now turn to the proof(s) of parts c) and d). Our first proof is purely functional analytic.
First proof of Theorem 3.1, c) and d).
Without loss of generality, $E:=\overline{\mathrm{span}}F(\Omega)$. Define the linear functional $T:E^{\prime}\to\mathbb{C}$ by
$T\mu:=\varphi(z\mapsto\langle F(z),\mu\rangle_{E,E^{\prime}})\qquad(\mu\in E^{\prime}).$
We claim that $T$ is induced by evaluation on a particular function $h\in E$, i.e.,
$\varphi(z\mapsto\langle F(z),\mu\rangle)=\langle h,\mu\rangle\qquad(\mu\in E^{\prime}).$
One then concludes $h(t)=\varphi(f_{t})$ for $t\in\Omega$ by specializing $\mu=\delta_{t}$.
It remains to prove the claim. Since $F$ is continuous, $E$ is separable and hence the $*{}^{*}$-topology on the unit ball of $E^{\prime}$ is metrizable. Since $\varphi$ is bp-continuous, $T$ is $*{}^{*}$ sequentially continuous on the unit ball of $E^{\prime}$, and hence $*{}^{*}$ continuous there. Applying a well-known theorem of Banach (see [17, Lemma 1.2] for an elegant proof) concludes the proof. ∎
Our second proof needs concepts from measure theory. Here, we suppose in addition that $\varphi$ is integration with respect to a measure. (Cf., however, Remark 3.2 below.)
Second Proof of Theorem 3.1, c) and d).
Let the functional $\varphi:\mathrm{H}^{\infty}(O)\to\mathbb{C}$ be given by integration against a complex Borel measure $\nu$. As before, we may suppose that $E=\overline{\mathrm{span}}(F(O))$. By a), $F:O\to E$ is holomorphic. Then $F$ is weakly holomorphic with values in a separable subspace. Hence, by Pettis’ measurability theorem, $F$ is strongly $|\nu|$-measurable. Since $F$ is bounded, it is Bochner integrable with respect to $|\nu|$, and
$h:=\int_{O}F(z)\,\mathrm{d}{\nu}\in E.$
Applying $\mu\in E^{\prime}$ yields
$\langle h,\mu\rangle=\int_{O}\langle F(z),\mu\rangle\,\mathrm{d}{\nu}=\varphi(z\mapsto\langle F(z),\mu\rangle).$
As before, specializing $\mu=\delta_{t}$ for $t\in\Omega$ we find $h(t)=\varphi(f_{t})$ and we are done. ∎
Remark 3.2.
The assumption that $\varphi$ is integration with respect to a measure is only apparently a restriction. Indeed, it turns out that each bp-continuous functional on $\mathrm{H}^{\infty}(O)$ is in fact given by integration against some finite measure, and even one that has a density with respect to Lebesgue measure. This follows from identifying the space of bp-continuous functionals on $\mathrm{H}^{\infty}(O)$ with the dual of $\mathrm{H}^{\infty}(O)$ with respect to the so-called mixed topology, see [6, Chap. V.1], in particular part 5) of Proposition 1.1 and Proposition 1.2. (Actually, Cooper [6] only treats the case $d=1$, but we would expect the same result for $d>1$.)
4. Holomorphic Families of Integrable Functions
In the present section we replace the set $\Omega$ by a measure space $(\Omega,\Sigma,\mu)$. The corresponding $\mathrm{L}_{p}$-space is denoted by $\mathrm{L}_{p}(\Omega)$, and $p,q$ are always dual exponents, i.e., $1\leq p,q\leq\infty$ with $\frac{1}{p}+\frac{1}{q}=1$.
Before we turn to the main results of this section, let us fix some auxiliary information.
@@ 4. Holomorphic Families of Integrable Functions
Then $|m(z,t)|\leq 1$ for all $(z,t)\in O\times\Omega$. Hence, $m$ meets the conditions of Theorem 3.1 with $E=\mathrm{BM}(X,\Sigma)$, the space of bounded $\Sigma$-measurable functions on $\Omega$. Given a bp-continuous functional $\varphi$ on $\mathrm{H}^{\infty}(O)$ it follows that
@@
The next is our second main theorem. Observe that assertion d) is an abstract Fubini-type result as in (1.1).
@@
$t\mapsto D_{z}^{\alpha}(a,t),\quad t\mapsto\lVert f_{t}\rVert_{\infty}\quad\text{and}\quad t\mapsto\varphi(f_{t})$
@@
where $\hskip-1.00006pt\left[\,F(z)\neq 0\,\right]=\{t\in\Omega\,\,|\,\,f(z,t)\neq 0\}$. The right-hand side is a $\sigma$-finite subset of $\Omega$. Hence, we may and shall from now on suppose that $\mu$ is $\sigma$-finite and fix measurable subsets $\Omega_{n}\subseteq\Omega$ with $\mu(\Omega_{n})<\infty$ and $\Omega_{n}\nearrow\Omega$.
@@
From part b) of Theorem  4.2 it follows that $D^{\alpha}F(a)g=D_{z}^{\alpha}f(a,\cdot)g$ $\mu$-almost everywhere. As $g\in\mathrm{L}_{1}(\Omega)$ is arbitrary here and $D_{z}^{\alpha}f(a,\cdot)$ is measurable, this implies that $D^{\alpha}F(a)=D_{z}^{\alpha}f(a,\cdot)$ $\mu$-almost everywhere. Hence, a) is proved.
Part c) of Theorem  4.2 implies that $t\mapsto\varphi(f_{t})g(t)$ is integrable with
@@
which is part c) halfway. Again by Theorem  4.2 d) applied with $h=\mathbf{1}$, yields d). Finally, it follows from d) and a standard application of the Hahn–Banach theorem that the function $t\mapsto\varphi(f_{t})$ (as an element of $\mathrm{L}_{\infty}(\Omega)$) is contained in the $\sigma(\mathrm{L}_{\infty},\mathrm{L}_{1})$-closure of $\mathrm{span}F(O)$. This concludes the proof. ∎
@@
1) For $d{=}1$ and $p{=}1$, some parts of Theorem 4.2 have been proved by Mattner in [15]. Mattner also has shown in [15, Counterexample 1] that if one replaces the first condition in Theorem 4.2 by the weaker one

### v2

@@ Abstract.
vector-valued, holomorphic functions of several variables, linearization theorem
@@ 1. Introduction
The paper is organized as follows. In Section 2 we recall classical results for vector-valued holomorphic functions. Next, in Section 3, we state a corollary of Mujica’s linearization theorem from [19], but with a new proof. In Section 4, we consider first the case of a closed subspace $E$ of the space $\mathrm{BM}(\Omega,\Sigma)$ of bounded measurable functions. The main result is then Theorem 4.3. It treats the case $E=\mathrm{L}_{p}(\Omega)$ for $1\leq p<\infty$. It is complemented by Theorem 4.4, which deals with $E=\mathrm{L}_{\infty}(\Omega)$.
@@
A subset $M\subseteq\ell^{\infty}(\Omega;E)$ is called bp-closed if it is closed under taking bp-limits of sequences in $M$.
Let $F\subseteq\ell^{\infty}(\Omega;E)$ be a bp-closed subspace. A linear mapping $\varphi:F\to X$ (where $X$ is any topological vector space) is called bp-continuous if
@@
We agree that it would be more accurate to speak of “sequentially bp-closed” sets and “sequentially bp-continuous” mappings. However, we decided to drop the word “sequentially” for the sake of brevity.
For any measurable space $(\Omega,\Sigma)$ we let $\mathrm{BM}(\Omega):=\mathrm{BM}(\Omega,\Sigma)$ be the Banach space of all bounded and measurable $\mathbb{C}$-valued functions, endowed with the supremum norm. This is a bp-closed subspace of $\ell^{\infty}(\Omega)$.
@@ 2. Vector-Valued Multivariate Holomorphic Functions
Here and in the following, $O\subseteq\mathbb{C}^{d}$ is a fixed open and not-empty set and $E$ is a complex Banach space. A function $f:O\to E$ is holomorphic if it is totally differentiable (= Fréchet-differentiable) with $\mathbb{C}$-linear derivative at each point of $O$ [14, Def. 147, p.68]. The following theorem is a useful characterization of holomorphy. It extends well-known results for multi-variable scalar-valued and one-variable vector-valued functions.
@@
We first show as in [16, proof of 1.3] that $f$ is continuous at $a$. Let $z=(z_{j})_{j}\in D$ and write
@@
is a well-defined $E$-valued holomorphic function on $D$. Indeed, $g$ is certainly holomorphic in each variable separately and each partial derivative is continuous. Hence one can apply [17, XIII, Thm. 7.1].
@@
1) For univariate functions, Theorem 2.1 is well-known, see [3, Appendix A]. For scalar-valued functions, the equivalence (i) $\Leftrightarrow$ (ii) is called Osgood’s lemma; the (a priori) stronger equivalence (i) $\Leftrightarrow$ (iii) is [16, Thm. 1.3].
2) Many books start from a different definition of holomorphy than ours and do not even mention Fréchet-differentiability. A noteworthy exception is [14]. See also [14, Theorem 160] for further characterizations of holomorphy.
3) Assertion (iii) involving a norming subset is due to Grothendieck [11], see also [15, p.139]. It implies (via the Hahn–Banach theorem) the following equivalences:
@@
    A function $f:O\to E$ is holomorphic if and only if it is weakly holomorphic (Dunford’s theorem, see [14, Thm. 148, p.68]).
@@
    A function $T:O\to\mathcal{L}(E;F)$ is holomorphic if and only if it is strongly holomorphic, if and only if for each $x\in E$ and each $x^{\prime}\in N$ from a norming subset $N\subseteq F^{\prime}$ the function $\langle F(\cdot)x,x^{\prime}\rangle$ is holomorphic.
4) Theorem 2.1 still holds if $N$ is merely a separating subset of $E^{\prime}$. This follows from the analogous result for univariate functions, due to Grosse-Erdmann [10]. Different proofs have been given by Arendt and Nikolski [4, Thm. 3.1] (see also [2]) and Grosse-Erdmann [9, Thm. 1]. For more general results in this direction see Frerick, Jordá and Wengeroth [8], in particular their Theorem 3.2.
@@
This follows from Hartogs’ theorem, which says that a scalar multi-variable function is already holomorphic if it is merely separately holomorphic [14, Thm. 153, p.69].
@@
We write $\mathrm{H}^{\infty}(O;E)$ for the space of all bounded holomorphic $E$-valued functions on $O$. If we endow $O$ with the Borel $\sigma$-algebra, $\mathrm{H}^{\infty}(O;E)$ becomes a (sequentially) bp-closed subspace of $\mathrm{BM}(O;E)$. From the Cauchy integral formula it follows that for fixed $a\in O$ and $\alpha\in\mathbb{N}_{0}^{d}$ the mapping
@@
3. The Linearization Theorem
The following theorem is a corollary of Mujica’s linearization theorem [19, Thm. 2.1], however with a different proof (see Remarks 3.2 below).
Theorem 3.1 (Linearization).
Let $O\subseteq\mathbb{C}^{d}$ be open, $E$ a Banach space and $f\in\mathrm{H}^{\infty}(O;E)$. Then for each bp-continuous functional $\varphi\in\mathrm{H}^{\infty}(O)^{\prime}$ there is a unique element $\varphi_{f}\in E$ such that
$\langle\varphi_{f},x^{\prime}\rangle=\langle x^{\prime}\circ f,\varphi\rangle\qquad(x^{\prime}\in E^{\prime}).$
Moreover, $\varphi_{f}\in\overline{\mathrm{span}}f(O)$.
Proof.
Uniqueness and the second assertion follow from the Hahn–Banach theorem. For existence we may suppose without loss of generality that $E=\overline{\mathrm{span}}f(O)$. As $f$ is continuous and $O$ is separable, so is $E$. Define the linear functional $\varphi_{f}:E^{\prime}\to\mathbb{C}$ by
$\varphi_{f}(x^{\prime}):=\varphi(z\mapsto(x^{\prime}\circ f)(z))\qquad(x^{\prime}\in E^{\prime}).$
We need to prove that $\varphi_{f}\in E$ under the natural embedding $E\hookrightarrow E^{\prime\prime}$. Since $E$ is separable, the weak-topology on the unit ball of $E^{\prime}$ is metrizable. Since $\varphi$ is bp-continuous, $\varphi_{f}$ is weakly sequentially continuous on the unit ball of $E^{\prime}$, and hence weakly continuous there. Then, by a well-known theorem of Banach (see [21, Lemma 1.2] for an elegant proof), it follows that $\varphi_{f}\in E$ as claimed. ∎
Remarks 3.2.
1) As it stands, Theorem 3.1 is a consequence of Mujica’s linearization theorem [19, Theorem 2.1]. There, Theorem 3.1 is stated for open $O\subseteq F$, where $F$ is any Banach space, and $\varphi\in\mathrm{G}^{\infty}(O)$. Here, $\mathrm{G}^{\infty}(O)$ is the space of all linear functionals which on bounded subsets of $\mathrm{H}^{\infty}(O)$ are continuous with respect to $\tau_{c}$, the topology of uniform convergence on compacts.
Now, since any open set $O\subseteq\mathbb{C}^{d}$ is locally compact and $\sigma$-compact, the topology $\tau_{c}$ is metrizable, and hence on bounded subsets of $\mathrm{H}^{\infty}(O)$, bp-continuity and $\tau_{c}$-continuity of a functional coincide.
2) Theorem 3.1 remains true (with essentially the same proof) under the following weaker hypotheses: $F$ is any Banach space, $O\subseteq F$ open, and $f\in\mathrm{H}^{\infty}(O;E)$ is separably-valued.
Under these hypotheses, a bp-continuous functional may not be $\tau_{c}$-continuous on bounded sets, and hence the result may then not be covered by Mujica’s theorem. (Unfortunately, we do not know of an example showing that this is indeed the case.)
3) Our proof of Theorem 3.1 is more elementary than Mujica’s from [19], as it avoids the bipolar theorem and is built just on Banach’s theorem (which has a much more elementary proof). Moreover, a slight modification of our argument also works in the more general setting of Mujica’s theorem: Namely, if $(x_{\alpha}^{\prime})_{\alpha}$ is a norm-bounded net and weakly-convergent to $x^{\prime}\in E^{\prime}$, then the net $(x_{\alpha}^{\prime}\circ f)_{\alpha}$ is $\tau_{c}$-convergent to $x^{\prime}\circ f$, by equicontinuity.
Using this argument leads to a simplification of the proof of Mujica’s theorem. However, it should be noted that the linearization result is only one (although essential) part of Mujica’s original theorem from [19].
The proof of Theorem 3.1 is primarily functional-analytic. It may be interesting to see that one can alternatively employ concepts from measure theory. Here, we suppose in addition that $\varphi$ is integration with respect to a measure. (Cf., however, Remark 3.3 below.)
Second Proof of Theorem 3.1.
Let the functional $\varphi:\mathrm{H}^{\infty}(O)\to\mathbb{C}$ be given by integration against a complex Borel measure $\nu$. Since $f$ is holomorphic, it is weakly holomorphic and has values in a separable subspace. Hence, by Pettis’ measurability theorem, $f$ is strongly $|\nu|$-measurable. Since $f$ is bounded, it is Bochner integrable with respect to $|\nu|$, and hence the Bochner integral
$h:=\int_{O}f\,\mathrm{d}{\nu}\in E$
exists. Applying $x^{\prime}\in E^{\prime}$ yields
$\langle h,x^{\prime}\rangle=\int_{O}\langle f(\cdot),x^{\prime}\rangle\,\mathrm{d}{\nu}=\varphi(z\mapsto\langle f(z),x^{\prime}\rangle).\qed$
Remark 3.3.
The assumption that the bp-continuous functional $\varphi\in\mathrm{H}^{\infty}(O)^{\prime}$ is integration with respect to a measure is only virtually restrictive. Indeed, it turns out that each bp-continuous functional on $\mathrm{H}^{\infty}(O)$ is already given by integration against some finite measure, and even one that has a density with respect to Lebesgue measure. This follows from identifying the space of bp-continuous functionals on $\mathrm{H}^{\infty}(O)$ with the dual of $\mathrm{H}^{\infty}(O)$ with respect to the so-called mixed topology, see [6, Chap. V.1], in particular part 5) of Proposition 1.1 and Proposition 1.2. (Actually, Cooper [6] only treats the case $d=1$, but we expect the same result for $d>1$.)
4. Holomorphic Families of Measurable and Integrable Functions
The Linearization Theorem 3.1 takes the form of an abstract Fubini-type theorem if $E$ is a space of functions on a set $\Omega$. We exploit this idea first for Banach spaces of bounded functions. Note that every space $E=\mathrm{BM}(\Omega,\Sigma)$ for some $\sigma$-algebra $\Sigma$ on a set $\Omega$ is a closed subspace of $\ell^{\infty}(\Omega)$.
@@ 4. Holomorphic Families of Measurable and Integrable Functions
Note that the second part of b) is of the type “differentiation under the integral” when one is inclined to interpret the application of $\mu$ as some kind of integral.
Proof of Theorem 4.1.
@@
are all holomorphic, $F$ is holomorphic by Theorem 2.1. It follows that $D^{\alpha}F$ takes values in $E$ and
@@
as scalar functions on $O$ for each $\mu\in E^{\prime}$ (Lemma 2.4). Specializing $\mu=\delta_{t}$ for $t\in\Omega$ yields $(D^{\alpha}F)(z)(t)=D^{\alpha}_{z}f(z,t)$, and hence b) is proved.
Next, we apply Theorem 3.1 to the mapping $F\in\mathrm{H}^{\infty}(O;E)$ and obtain, for any bp-continuous functional $\varphi\in\mathrm{H}^{\infty}(O)^{\prime}$ an element $\varphi_{F}\in\overline{\mathrm{span}}(F(O))$ with
$\langle\varphi_{F},\mu\rangle=\varphi(z\mapsto\langle F(z),\mu\rangle)\qquad(\mu\in E^{\prime}).$
By specializing $\mu=\delta_{t}$ for $t\in\Omega$, we find $\varphi_{F}=(t\mapsto\varphi(f_{t}))$. This concludes the proof of c) and d). ∎
Next, we replace the set $\Omega$ by a measure space $(\Omega,\Sigma,\mu)$. The corresponding $\mathrm{L}_{p}$-space is denoted by $\mathrm{L}_{p}(\Omega)$, and $p,q$ are always dual exponents, i.e., $1\leq p,q\leq\infty$ with $\frac{1}{p}+\frac{1}{q}=1$.
Before we turn to the main results, let us fix some auxiliary information.
@@
Then $|m(z,t)|\leq 1$ for all $(z,t)\in O\times\Omega$. Hence, $m$ meets the conditions of Theorem 4.1 with $E=\mathrm{BM}(\Omega,\Sigma)$, the space of bounded $\Sigma$-measurable functions on $\Omega$. Given a bp-continuous functional $\varphi$ on $\mathrm{H}^{\infty}(O)$ it follows that
@@
The next is our main theorem. Observe that assertion d) is an abstract Fubini-type result as in (1.1).
@@
$t\mapsto D_{z}^{\alpha}f(a,t),\quad t\mapsto\lVert f_{t}\rVert_{\infty}\quad\text{and}\quad t\mapsto\varphi(f_{t})$
@@
where $\hskip-1.00006pt\left[\,F(z)\neq 0\,\right]=\{t\in\Omega\,\,|\,\,f(z,t)\neq 0\}$. The right-hand side is a $\sigma$-finite subset of $\Omega$. Hence, we may suppose that $\mu$ is $\sigma$-finite. Accordingly, we fix measurable subsets $\Omega_{n}\subseteq\Omega$ with $\mu(\Omega_{n})<\infty$ and $\Omega_{n}\nearrow\Omega$.
@@
From part b) of Theorem 4.3 it follows that $D^{\alpha}F(a)g=D_{z}^{\alpha}f(a,\cdot)g$ $\mu$-almost everywhere. As $g\in\mathrm{L}_{1}(\Omega)$ is arbitrary here and $D_{z}^{\alpha}f(a,\cdot)$ is measurable, this implies that $D^{\alpha}F(a)=D_{z}^{\alpha}f(a,\cdot)$ $\mu$-almost everywhere. Hence, a) is proved.
Part c) of Theorem 4.3 implies that $t\mapsto\varphi(f_{t})g(t)$ is integrable with
@@
which is part c) halfway. Again by Theorem 4.3 d) applied with $h=\mathbf{1}$, yields d). Finally, it follows from d) and a standard application of the Hahn–Banach theorem that the function $t\mapsto\varphi(f_{t})$ (as an element of $\mathrm{L}_{\infty}(\Omega)$) is contained in the $\sigma(\mathrm{L}_{\infty},\mathrm{L}_{1})$-closure of $\mathrm{span}F(O)$. This concludes the proof. ∎
@@
1) For $d{=}1$ and $p{=}1$, some parts of Theorem 4.3 have been proved by Mattner in [18]. Mattner also has shown in [18, Counterexample 1] that if one replaces the first condition in Theorem 4.3 by the weaker one